OpenStudy (anonymous):

Life of Light Bulbs A certain type of light bulb has an average life of 500 hours, with a standard deviation of 100 hours. The length of life of the bulb can be closely approximated by a normal curve. An amusement park buys and installs 10,000 such bulbs. Find the total number that can be expected to last for each period of time. Between 680 and 780 hours.

OpenStudy (anonymous):

z=680-500/100=1.8 z=780-500/100=2.8

OpenStudy (phi):

yes, looks good so far. use a Z-table to find the % area between those Z scores, and multiply by 10,000 to get the expected number that last between the given limits.

OpenStudy (anonymous):

idk how to use the ztable it confuses me

OpenStudy (phi):

From http://en.wikipedia.org/wiki/Standard_normal_table#Cumulative_table look for 1.8 on the left side. use the first column (which is 1.80), to get the area up to +1.8 std dev. what do you get ?

OpenStudy (anonymous):

0.9641

OpenStudy (phi):

now do the same for 2.8

OpenStudy (anonymous):

and .9974

OpenStudy (phi):

subtract 0.9974- 0.9641

OpenStudy (phi):

multiply by 10,000 to get the expected number that last between the given limits.

OpenStudy (anonymous):

.0333 333

OpenStudy (phi):

notice that if you wanted the z score for 1.81 you would go over 1 column and read 0.9649

OpenStudy (anonymous):

ok, what do i do if i a ztable isnt furnished for me for a test

OpenStudy (phi):

They may expect you to memorize the area between the mean and +1 std dev, at 2 std, and 3 std dev (3 numbers) But they won't ask questions like this unless they give you the z table.

OpenStudy (anonymous):

ok. i thank you for helping me